Find the equation of the tangent line to each curve when x has a given value.



1. f(x)= x²+2 as x= 2


2.f(x)= 3x²+1 as x=1



Find the slope of the tangent line of the given function f below.



1.f(x)=2x+5


2.f(x)=x²-1

1
Expert's answer
2022-03-07T17:21:05-0500

equation of the tangent line at the point "x_0" is:

"y-y_0=y'(x_0)(x-x_0)"

1. f(2) = 6

"f'(x)=2x\implies f'(2)=4"

so, the equation of the given line is

"y-6=4(x-2)"


2. f(1) = 4

"f'(x)=6x\implies f'(1)=6"

so, the equation of the given line is

"y-4=6(x-1)"


The slope of the tangent line to a function f(x) is f'(x), so

1.f'(x) = 2


2.f'(x) = 2x